

********************************************************************
UARS WEEKLY STATUS REPORT - 21 November 1997
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
This week also performed a special solar pointing experiment to help us
characterize the SSPP's stellar pointing performance. This was similar to the
exeriment we did last July (see weekly report 08/01/97). Reults showed a
pointing error near 30 arc-seconds and good agreement with a nominal solar
track. This give us confidence in the ability of the SSPP to accurately
track a stellar target.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems.
***********
PEM
***********
PEM AXIS were active from November 9 through November 11. During this
time, PEM produced energy deposition profiles due to electrons inverted from
X-ray data.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI was turned off as scheduled on the 19th of November, 1997. Our next =
scheduled on-time will be December 8. Since next week a holiday, we wish =
everyone involved a Happy Thanksgiving.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
UARS CDHF Status Week Ending Nov 21, 1997
-----------------------------------------
UARS Level 0, NRT and Quicklook file generation software was successfully
re-hosted to the new hardware platform on November 15.
The CDHF has received a total of 346 CD-ROM requests as of 21 November 97.
Five new CD-ROM orders were received and shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) Prod (1350R-590C=760L)
The CDHF reingested level 0 for UARS day 2166 due to better data quality.
The CDHF started receiving UKMO data on 14-November-1997. UKMO is now catching
up with missing days.
Pending CDHF software upgrades include IDL 5.0 and OpenIngres 2.0.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 November 1997 (Orbit 33653, GMT Day 311) through
14 November 1997 (Orbit 33757, GMT Day 318).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 313/0013 33668 AXIS On
HALOE 315/0022 33698 On
HALOE 315/0202 33700 Low power
HALOE 315/1601 33708 On
PEM 315/1815 33710 AXIS Off
HRDI 316/0029 33714 On
HRDI 316/0206 33715 Off
HRDI 317/0029 33729 On
HRDI 317/0212 33730 Off
HRDI 317/0623 33732 On
The beta angle is now increasing toward a maximum angle of 78.1 degrees
on 27 November 1997 (DOY 331).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery
1 remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery
2 and Battery 3 of 1.4 degrees C. Additional battery performance data
for each battery on 08 November 1997 and 14 November 1997 are provided
in Appendix D.
The clock error ranged from -11.2 to +9.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 19
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HRDI 06
% HALOE 03
% SUSIM 01
% SOLSTICE 01
% PEM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
TRANSPONDER FREQUENCIES:
Transponder A: 97/301 - 2287.497086
Transponder B: 97/302 - 2287.499098
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 97/317 show the observatory data loss* to
be 20 hours, 41 minutes, 00.148 seconds (an increase of 4.192 seconds).
This is a 0.0384 percent data loss which equals a 99.9616 percent data
capture for the mission.
The increase of 4.192 seconds occurred on Day 317 and was due to a hit
in the data.
*NOTE
The total data loss listed above includes a correction for a loss on Day
228 that was recovered. Part of the data loss on Day 228 was attributed
to hits in the data but was really due to an in-house line problem. The
Level-0 has been recreated to include the missing data and forwarded to
CDHF.
UARS POCC database version 064 was officially released on 07 November
1997. The database incorporated minor modifications to some of the
partial system tables (p-tables) and discrete state definitions. The
database was tested by the FOT before being released to operations.
Flight Dynamics Facility (FDF) continues the generation process for a
new Star Catalog to fix the invalid precession rate for star 48000001.
All previous star catalogs were generated by FDF on the IBM system.
This will be the first catalog created using the HP workstation, and the
procedure to generate the Star Catalog is still being ironed out.
The FOT is still working out the details of participating in the IP
return-link telemetry test that the network is conducting through the
end of November 1997. There is still some concern regarding the
integrity of the IP conversion devices that UARS will be utilizing.
The FOT received a request to perform a special solar tracking
experiment from SOLSTICE operations this week. The experiment (which
was performed earlier this year) will involve commanding the SSPP to
point to a particular Right Ascension and Declination, instead of using
the Solar ephemeris to track the Sun. All of the pointing times and
angles will be calculated by SOLSTICE operations and confirmed by the
FOT.
No testing of the new CMS software release (R7B1V2) occurred this week.
Investigation of the stored command verification problem - time to
complete - associated with Applications Processor (AP) software release
10.08 continued. The cause has been identified, but rectifying the
problem continues.
OTHER SIGNIFICANT EVENTS
HALOE FOT submitted a contingency plan to save power and avoid a full
shut down during reduced power situations. The HALOE FOT maintains that
this contingency plan removes 25 Watts from the bus. The contingency
plan works by turning off the HALOE heater/motor bus and leaving their
Quiet Bus (QB) and Pulse Bus (PB) on. The FOT utilized this mode on Day
315 during the pass after HALOE was completely powered up to full power
thereby avoided shut down. The following morning HALOE was fully
powered up.
Monitoring of the magnetic field model performance continues.
FOT continues to develop an updated Power Profile for UARS due to the
current status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering
effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not
been scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 17 December
1997 at 311/132708.
The Replacement Level Zero Processor (LZPR) will be elevated to
operational status on 15 November 1997. The replacement procedure will
result in a planned 10-hour downtime during which no Level Zero Data
will be processed - no Near Realtime Data (NRT) or FOT plots.
The UARS orbit altitude is dropping slowly and a drag makeup orbit
adjust will probably be required in mid-1998.
FACILITIES
UARS Off-line Engineering moved from Room E137 on the first floor of
Building 32 to Room W235 on the second floor adjacent to the newly
refurbished UARS/TOMS control center on 10 November 1997. Telephone
numbers are unchanged.
The UARS and TOMS control centers were relocated to the newly
refurbished unified control center on 11 November 1997. The new control
center with new consoles is located in the original location for UARS in
Room C221 of Building 32.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 14 November 1997
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
33668 313/0013 97-555 PEM Power ON Axis
33682 313/2237 97-556 SUSIM Commands
33698 315/0022 97-557 HALOE Power ON
33700 315/0202 97-558 HALOE HALOE To Low Power Mode
33708 315/1601 97-559 HALOE Returned to Normal Ops
33710 315/1815 97-560 PEM Power OFF AXIS
33714 316/0029 97-561 HRDI Power ON
33715 316/0206 97-562 HRDI Power OFF
33719 316/0950 97-563 OBC Clk Adj fr 0.00 to 20.6
msec/day
33729 317/0029 97-564 HRDI Power ON
33730 317/0212 97-565 HRDI Power OFF
33732 317/0623 97-566 HRDI Reduce HTR delay to 400
amp-min
33732 317/0623 97-567 HRDI Power ON
33743 317/2253 97-568 OBC Clk Adj fr 20.6 to 0.00
msec/day
33758 318/2337 97-568 SOLSTICE Reboot Soltice
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 November 1997
The following Out-of-Limit (ALARM) conditions were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
CDUNREG28 Y-LO CDH +28V UNREG BUG 33702
HAH28VA R-LO +28 HTR/MTR PS VOLT 33700-8
HRPSEPBIMON R-LO HRDI PULSE BUS CURRENT MONITOR 33730
HRPTSPBIMON Y-HI HRDI TELESCOPE PB CURRENT MON. 33734
NBRECBI Y-HI NBTR B Current 33657,8,716
OBSSOC2 Y-LO Battery 2 Current SOC Continuous
OBSVLBEON R-LO Load Bus Volts at EON Continuous
OBSBATVEON2 R-LO Battery 2 voltage at EON Continuous
OBSBATVEON3 R-LO Battery 3 voltage at EON Continuous
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 33656-
60,4,73,88,717
APPENDIX B (cont)
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 14 November 1997
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
PWLBV Y-LO Current Load Bus Volts 33687,702,17,52
PWBAT2V Y-LO Battery 2 Current Voltage 33702,52
PWBAT3V Y-LO Battery 3 Current Voltage 33687,702,17,52
PWLBV R-LO Current Load Bus Voltage 33657,8,72
PWBAT2V R-LO Current Batt 2 Voltage 33657,8,72
PWBAT3V R-LO Current Batt 3 Voltage 33657,8,72
SOSGTEVLTG R-LO SOLSTICE G_Tube Voltage 33756
SORGTERVE R-LO SOLSTICE G_Tube R_Voltage 33756
WIEUTEMP Y-LO WINDII EU Temperature 33654-113
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 14 November 1997
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
33719 316/09:56 (12 Nov 97) from 00.0 to 20.6 msec/day
33742 317/22:56 (13 Nov 97) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 14 November 1997
08 Nov 1997 - GMT Day 312 Beta = 4.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.03/-3.89 +0.0/ +0.0 18.7 0.000 0.0
2 +33.6/ -5.6 +7.17/+5.85 +30.8/-15.6 23.8 1.020 24.2
3 +44.8/ -11.2 +5.79/+4.51 +29.8/-15.2 23.8 1.024 23.2
14 Nov 1997 - GMT Day 318 Beta = 30.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.94/-3.68 +0.0/ +0.0 18.7 0.000 0.0
2 +22.4/ -33.6 +7.74/+6.44 +31.0/-16.4 24.2 1.019 24.8
3 +33.6/ -22.4 +6.44/+5.15 +29.7/-16.1 24.0 1.021 23.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 14 November 1997
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
97-119 FOT 33698 315/0022 14170 Generic Late Acquisition
97-119 FOT 33708 315/1601 14170 Generic Late Acquisition
97-121 SOLSTICE 33757 318/2137 N/A Solstice Crash
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period